### v1

# Athena: Enhancing Multimodal Reasoning with Data-efficient Process Reward Models

Abstract
We present Athena-PRM, a multimodal process reward model (PRM) designed to evaluate the reward score for each step in solving complex reasoning problems. Developing high-performance PRMs typically demands significant time and financial investment, primarily due to the necessity for step-level annotations of reasoning steps. Conventional automated labeling methods, such as Monte Carlo estimation, often produce noisy labels and incur substantial computational costs. To efficiently generate high-quality process-labeled data, we propose leveraging prediction consistency between weak and strong completers as a criterion for identifying reliable process labels. Remarkably, Athena-PRM demonstrates outstanding effectiveness across various scenarios and benchmarks with just 5,000 samples. Furthermore, we also develop two effective strategies to improve the performance of PRMs: ORM initialization and up-sampling for negative data. We validate our approach in three specific scenarios: verification for test time scaling, direct evaluation of reasoning step correctness, and reward ranked fine-tuning. Our Athena-PRM consistently achieves superior performance across multiple benchmarks and scenarios. Notably, when using Qwen2.5-VL-7B as the policy model, Athena-PRM enhances performance by 10.2 points on WeMath and 7.1 points on MathVista for test time scaling. Furthermore, Athena-PRM sets the state-of-the-art (SoTA) results in VisualProcessBench and outperforms the previous SoTA by 3.9 F1-score, showcasing its robust capability to accurately assess the correctness of the reasoning step. Additionally, utilizing Athena-PRM as the reward model, we develop Athena-7B with reward ranked fine-tuning and outperforms baseline with a significant margin on five benchmarks.
1 Introduction
In recent years, Large Language Models (LLMs) [5, 1, 52, 30, 61] have achieved remarkable success in natural language processing. Building on this, Multimodal Large Language Models (MLLMs) [31, 26, 3, 8] have made significant strides in various vision-language tasks, such as visual question answering [2, 39] and chart understanding [38]. Despite their impressive performance, solving complex tasks involving mathematical and multi-step reasoning remains challenging.
To enhance reasoning capabilities, several approaches have been explored, including fine-tuning on long chain-of-thought (CoT) data [59, 40], offline preference optimization [56, 43], and online reinforcement learning [47, 17, 21, 51]. Another promising avenue is test time scaling (TTS), which involves generating multiple responses from a policy model and selecting the most consistent answer [58] or the solution with the highest reward using reward models [9, 28, 53, 50, 46, 55].
Reward models for TTS mainly include two types: outcome reward models (ORMs) [9] and process reward models (PRMs) [28]. ORMs evaluate the reward score for a given question and its solution, whereas PRMs provide reward scores for each intermediate reasoning step, offering fine-grained feedback. PRMs typically deliver superior performance and better out-of-distribution generalization [28]. However, obtaining high-quality data with process labels poses significant challenges. PRM800K [28] involves collecting 800K labeled steps through human annotations, which is time-consuming and requires skilled annotators, particularly for complex multi-step reasoning and mathematical tasks. Math-Shepherd [55] proposes an automated labeling method using Monte Carlo (MC) estimation. It defines the quality of an intermediate step as its potential to reach the correct answer. MC estimation for step-level labels is typically performed by sampling numerous reasoning trajectories using a language model, referred to as the completer, to estimate the probability of arriving at the correct answer. However, this approach involves significant computational overhead. Besides, another shortcoming of MC-based estimation methods is that estimated step-level labels are inevitably noisy.
We aim to address the above challenges: reducing computational costs and mitigating noisy step labels. We first find that the accuracy of step labels estimated by MC-based methods is influenced by the reasoning capability of the completer. A strong completer can arrive at the correct answer despite incorrect intermediate steps as shown in Figure 1, whereas a weak completer may struggle even with correct intermediate steps. The estimation of MC-based methods is biased toward the completers we use. However, the correctness of the intermediate step should not depend on the completer. Based on this insight, we propose using both weak and strong completers to estimate step labels, retaining only those steps where labels generated by both completers are consistent to remove the bias caused by completers. This approach improves the quality of the step label. Empirical results demonstrate that a small set of high-quality labels ($\approx$ 5K) achieves impressive performance compared to large-scale data ($\approx$ 300K) labeled by vanilla MC-based estimation [55]. In addition, our approach requires only 1/45 of the GPU hours for data synthesis and 1/60 of the GPU hours for reward model training, significantly lowering computational costs.
Figure 1: Illustration of different completers under the same question and solution steps. Even if given wrong intermediate steps, the strong completer stills reach the final answer while the weak completer fails. We omitted some intermediate steps in the figure for simplicity.
After acquiring high-quality datasets, we explore two effective strategies for training PRMs: initialization from ORMs and up-sampling data with negative labels. PRMs are typically fine-tuned from pre-trained foundation models, such as LLaMA-3.1-8B-Instruct [16] in RLHF-workflow [13]. Previous studies indicate that ORMs possess some ability to assess intermediate step correctness through confidence variation [33] or parameterization [10, 63]. Inspired by this, we find that initializing PRMs from ORMs significantly boosts performance, as ORMs trained on large-scale response-level data serve as pre-training with weak supervision. And PRMs can be considered fine-tuning on high-quality, fine-grained step-level data. Additionally, we show that label imbalance exists in process-labeled solutions and we propose to up-sample data with negative step labels to address this.
Building on these methodologies, we develop our outcome reward model Athena-ORM and process reward model Athena-PRM. Leveraging Athena-PRM, we introduce Athena-7B with reward ranked fine-tuning. We validate our approach across three scenarios: 1) test time scaling  [50]: where Athena-PRM ranks multiple outputs generated by policies under the Best-of-N evaluation; 2) direct evaluation of the correctness of reasoning steps using Athena-PRM; 3) reward ranked fine-tuning [12]: where Athena-PRM ranks outputs sampled from the current policy, utilizing the highest reward response for policy model fine-tuning. In the TTS scenario, we evaluate Athena-PRM on seven multimodal math and reasoning benchmarks with three different policy models ranging from 7B to 72B, demonstrating significant improvements in reasoning abilities. For instance, on the WeMath benchmark [44], Athena-PRM enhances the zero-shot baseline by 10.2 points using Qwen2.5-VL-7B [3] as the policy model. Furthermore, Athena-PRM excels in text-only benchmarks, achieving an 8.9 points improvement on a challenging math benchmark [18] with Mistral-8B. To assess its ability to judge intermediate reasoning step correctness, Athena-PRM is evaluated on the VisualProcessBench [57], showcasing strong performance and outperforming VisualPRM-8B [57], an open-source multimodal PRM, and proprietary models as judge. In the reward ranked fine-tuning scenario, our fine-tuned model Athena-7B, based on Qwen2.5-VL-7B [3], significantly enhances the policy model’s reasoning capabilities across seven math and reasoning benchmarks.
Our contributions are summarized as follows:
  • $\bullet$
    We propose using prediction consistency between weak and strong completers to filter noisy process labels, enhancing the quality of the automated-labeled process data. The high-quality process labeled data shows surprising data efficiency for training PRMs.
  • $\bullet$
    We introduce two effective strategies to improve the performance of PRMs: ORM initialization and negative data up-sampling.
  • $\bullet$
    We develop our reward models Athena-ORM and Athena-PRM, and evaluate them under the Best-of-N setting on nine math and reasoning benchmarks across various model sizes and families, demonstrating the effectiveness of our approach. Moreover, Athena-PRM achieves state-of-the-art (SoTA) performance in assessing the correctness of intermediate steps directly on the VisualProcessBench [57].
  • $\bullet$
    Leveraging Athena-PRM, we introduce Athena-7B, a MLLM with exceptional capabilities fine-tuned using reward ranked fine-tuning. Extensive experiments highlight the superiority of Athena-7B across diverse benchmarks.
2 Method
In this section, we first introduce basic concepts, including ORMs and PRMs in Sec. 2.1. Next, we present our method for constructing a high-quality PRM training set, along with practical training strategies to improve PRM performance in Sec. 2.2. The data collection process for training is detailed in Sec. 2.3. Finally, we discuss three application scenarios of PRMs in Sec. 2.4.
2.1 Reward Models for Mathematical Problem
ORMs.  Given a mathematical problem $x$ with golden answer $y^{*}$ and its solution $a$ generated by policy $\pi$, i.e., $a\sim\pi\left(\cdot\mid x\right)$, ORMs assign a reward $r\left(x,y,y^{*}\right)$ to the final answer $y$ of $a$ to reflect the correctness of the solution. To train ORMs, we take inputs as $(x,y)$ and predict correctness $\delta(y,y^{*})$ of prediction with following loss function:
$\mathcal{L}_{ORM}=\delta\left(y,y^{*}\right)\cdot\log r\left(x,y,y^{*}\right)+\left(1-\delta\left(y,y^{*}\right)\right)\cdot\log\left(1-r\left(x,y,y^{*}\right)\right),$ (1)
where $\delta\left(y,y^{*}\right)=1$ if $y=y^{*}$, otherwise $\delta\left(y,y^{*}\right)=0$.
[…]
$\mathcal{L}_{PRM}=\sum_{i=1}^{K}\delta_{i}\cdot\log r\left(x,a_{i}\right)+\left(1-\delta_{i}\right)\cdot\log\left(1-r\left(x,a_{i}\right)\right),$ (2)
where $\delta_{i}=1$ if the step $a_{i}$ is correct, otherwise $\delta_{i}=0$.
2.2 Automated Data Scaling of PRMs
Compared with ORMs, PRMs usually achieve better performance and exhibit stronger out-of-distribution generalization [28]. However, collecting high-quality data with process labels is challenging. Human annotation provides accurate supervision signals [28], but it requires expensive human labeling and difficult to scale up. Math-Shepherd [55] proposes to use a Monte Carlo (MC) sampling method to estimate the correctness of each step without human supervision. Specifically, to estimate the correctness of step $a_{i}$, we use a completer $\phi$ to finalize the reasoning process from step $a_{i}$, and get the final answer $y$. We repeat sampling $T$ times and get corresponding answers $\mathcal{Y}=\{y_{j}\}_{j=1}^{T}$. There are two ways to estimate the correctness of $a_{i}$: soft estimation and hard estimation.
For soft estimation, we assume that the frequency of getting the correct answer could dictate the quality of a step:
$\delta_{i}^{soft}=\frac{\sum_{j=1}^{T}\mathbb{I}(y_{j}=y^{*})}{T}.$ (3)
For hard estimation, we assume that the correct answer is right when the step could reach the answer in $T$ samples:
$\delta_{i}^{hard}=\begin{cases}1&\text{if }\exists y_{j}\in\mathcal{Y},y_{j}=y^{*},\\ 0&\text{otherwise.}\end{cases}$ (4)
In this paper, we mainly discuss the hard estimation for PRMs because it allows us to use standard language modeling loss to train PRMs and eliminates the adjustments of the training pipeline.
MC-based estimation methods provide an automated and scalable labeling strategy for intermediate reasoning steps. However, automatically labeled data often contain incorrect labels, and MC-based methods typically incur substantial computational costs due to extensive sampling requirements. For instance, when processing a solution $a$ with $K$ steps, the completer $\phi$ must generate $T\times K$ solutions, which produces $T\times K$ times computation cost compared with ORMs.
To enhance the accuracy of the labeling process and reduce the computational burden of data synthesis, our central approach is to utilize a smaller set of high-quality data with process labels. We improve the quality of automatically labeled process labels by introducing consistency filtering between weak and strong completers. The rationale is straightforward: the results of hard estimation depend on the number of samples $T$, the difficulty of the problem $x$, and the ability of the completer $\phi$. A strong completer can arrive at the correct answer even when provided with incorrect steps, as shown in Figure 1, whereas a weak completer struggles to find the final answer even when given correct steps. We aim to enhance the quality of process labels through prediction consistency between weak and strong completers. Specifically, we only use steps whose assigned labels are consistent between different completers.
To validate our method, we conducted a simple experiment using the PRM800K dataset [28]. We employed Mistral-7B-Instruct [22] as the weak completer ($\phi_{w}$) and Qwen2.5-72B-Instruct [61] as the strong completer ($\phi_{s}$), setting the number of samples $T=8$ for MC-based estimation. We randomly sampled 50 queries and reported the accuracy of estimated process labels:
Weak Completer $\phi_{w}$ Strong Completer $\phi_{s}$ Consistency Filter
Accuracy (%) 78.2 83.4 94.1
Our experiments show that label quality improves significantly after applying the consistency filter. Empirically, we demonstrate that a small number of high-quality labels is sufficient to achieve superior performance, thereby reducing computational costs compared to the vanilla MC-based estimation [55]. Results of more strategies and combination of completers are provided in Appendix A.1.
Additional strategies for PRMs. We offer two strategies for enhancing the performance of PRMs: initialization from ORMs and negative data up-sampling. First, initializing PRMs with ORMs trained on large-scale sample-level annotated data improves performance. [^1] It is shown that initialization from ORMs will improve performance with different datasets, see verification in Table 5. Previous work indicates that ORMs possess a certain capacity to assess the correctness of intermediate steps [10, 33, 63]. We empirically show that PRMs benefit from a few high-quality examples when initialized with ORMs. Large-scale sample-level annotation data provides weaker supervision, with outcome supervision acting as a simplified form of process supervision. Training ORMs on large-scale sample-level annotations serves as a “pre-training”, while training PRMs from pre-trained ORMs acts as“fine-tuning” with high-quality process-level annotated data. Secondly, our findings reveal that label imbalance is prevalent in most process-labeled datasets, such as PRM800K [28] and our synthetic data, where correct steps are more common than errors. We provide a detailed distribution of process labels in Table 9. Empirically, slightly up-sampling data with negative labels improves performance with minimal additional computation.
2.3 Data Collection
To construct a diverse and high-quality dataset, we first gather data from various public multimodal and text-only datasets. For multimodal datasets, we collect from MathV360k [48], UniGeo [7], Geometry3k [34], CLEVR-Math [29], ScienceQA [35], GeomVerse [24], GeoQA-plus [6], and DocVQA [39]. For text-only datasets, we utilize GSM8K [9], MATH [18], and NuminaMath-1.5.222https://huggingface.co/datasets/AI-MO/NuminaMath-1.5 If the original dataset includes different splits, we use only the training portion to prevent test set contamination. We remove all judgment questions and convert multiple-choice questions to open-ended ones to deter policy models from “guessing” answers. Then, for each query, we sample 8 responses from the policy model. To ensure dataset diversity, we employ multiple models as policies, including InternVL2.5-8B/78B [8], Qwen2.5-VL-7/72B-Instruct [3], LLaVA-OneVision-7B/72B [26] for multimodal datasets, and Qwen2.5-7/72B-Instruct [61] for text-only datasets.
Next, we apply carefully designed filtering rules to all responses. We exclude responses with too many or too few tokens and eliminate those exhibiting repetitive patterns. We employ $n$-gram deduplication to remove similar responses and enhance diversity. Ultimately, we obtain approximately 600K queries with corresponding responses. We parse answers from responses and assign correctness labels of 0 or 1 to each query-response pair. We refer to this dataset as Athena-600K and employ it to train our outcome reward model Athena-ORM.
For training PRMs, we generate process labels using the proposed method introduced in Sec. 2.2. We use Qwen2.5-VL-7B [3] as the weak completer ($\phi_{w}$) and Qwen2.5-VL-72B [3] as the strong completer ($\phi_{s}$). Finally, we collect approximately 5K samples for training our process reward model Athena-PRM. We also compare with PRMs trained using approximately 300K samples estimated by the vanilla MC-based method outlined in [55]. We denote these datasets as Athena-5K and MC-300K, respectively, and compare their performance and computational cost in Sec. 3.4.
2.4 Application
Following reward model training, we assess the performance of ORMs and PRMs in three scenarios: test time scaling [50], direct judgment, and reward ranked fine-tuning [12].
Verification for test time scaling.  We adopt a Best-of-N evaluation paradigm. Specifically, given a problem $x$ in the test set, we sample $N$ solutions from the policy $\pi$. All solutions are scored using a reward model, and we choose the solution with the highest score. For ORMs, we directly use the outputs from ORMs as the reward of solutions. For PRMs, the minimum reward across all steps is used to rank all solutions. [^3] We test different choices, including the reward from the last step, the minimum reward across all steps, or the product of the rewards from all steps, as the reward for solutions in Appendix A.3.
Direct judgment for reasoning steps.  Besides verification at test time under the Best-of-N setting, Athena-PRM can also be used to directly identify erroneous steps in the mathematical reasoning process. Given a solution $a$ with $K$ steps as input, Athena-PRM outputs the correctness score $\{\delta_{i}\}_{i=1}^{K}$ for each step $\{a_{i}\}_{i=1}^{K}$ in a single forward pass.
Response ranking for reward ranked fine-tuning.  We explore the use of PRMs for data synthesis in reward ranked fine-tuning [12, 64, 49, 66]. Using the current policy $\pi$, we generate $M=8$ solutions for an input problem $x$. Subsequently, we filter out solutions with incorrect answers and apply deduplication to remove highly similar responses, enhancing the diversity of the remaining solutions. We retain queries where 2 to 6 out of 8 responses are correct, excluding those with too few or too many correct answers, as such cases are either too easy or too challenging for the current policy $\pi$. Following this filtering step, we use Athena-PRM to score all solutions for each query and select the solution with the highest reward as the corresponding label. The policy $\pi$ is fine-tuned using the synthetic data generated as described above.
3 Experiments
Table 1: Results on seven multimodal reasoning benchmarks under Best-of-N (N=8) evaluation. $\dagger$ denotes that the results are from [57]. We mark our results and highlight the best result.
WeMath MathVista MathVision MathVerse DynaMath MMMU LogicVista
Qwen2.5-VL-7B [3] 36.2 68.1 25.4 41.1 21.8 58.0 47.9
Self-consistency [58] 44.7 71.6 28.6 43.7 22.9 60.1 49.5
[…]
Athena-ORM 45.1 72.8 29.8 44.1 23.1 62.7 51.3
Athena-PRM 46.4 75.2 32.5 46.3 23.4 63.8 53.0
InternVL2.5-8B [8] 23.5 64.5 17.0 22.8 9.4 56.2 36.0
Self-consistency [58] 28.4 66.1 21.1 24.7 13.8 57.8 40.2
[…]
Athena-ORM 28.6 66.9 22.0 25.8 15.2 59.1 40.8
Athena-PRM 30.1 71.4 23.4 26.1 18.7 60.3 44.4
Qwen2.5-VL-72B [3] 49.1 74.2 39.3 47.3 35.9 70.2 55.7
Self-consistency [58] 54.8 77.0 43.1 50.8 37.6 71.1 59.6
Athena-ORM 55.6 77.8 43.0 51.2 39.6 72.3 60.1
Athena-PRM 58.7 79.1 44.8 54.6 42.5 75.8 60.9
Table 2: Results on text-only math benchmarks under Best-of-N (N=8) evaluation. $\dagger$ denotes that the results are from [57]. We mark our results and highlight the best result.
(a) Qwen2.5-7B
GSM8K MATH
Qwen2.5-7B [61] 91.6 75.5
Self-consistency [58] 93.4 79.9
[…]
Athena-ORM 94.0 81.1
Athena-PRM 94.8 82.4
(b) Ministral-8B
GSM8K MATH
Ministral-8B 85.6 54.5
Self-consistency [58] 90.7 62.0
Athena-ORM 91.4 63.2
Athena-PRM 93.1 65.4
(c) Qwen2.5-72B
GSM8K MATH
Qwen2.5-72B [61] 95.8 83.1
Self-consistency [58] 96.0 86.0
[…]
Athena-ORM 96.0 86.2
Athena-PRM 97.0 87.4
To comprehensively evaluate the proposed method, we assess Athena-ORM and Athena-PRM under the Best-of-N setting on seven multimodal benchmarks and two text-only benchmarks in Sec. 3.1. Furthermore, we evaluate Athena-PRM on VisualProcessBench, comparing its performance with MLLMs as judges and a recent open-source multimodal PRM, VisualPRM-8B [57], as described in Sec. 3.2. We assess the capability of Athena-7B on multimodal math and reasoning tasks in Sec. 3.3. To validate the effectiveness of our designs for Athena-PRM, an in-depth analysis is provided in Sec. 3.4.
3.1 Best-of-N Evaluation
Benchmarks. We evaluate Athena-ORM and Athena-PRM across multiple multimodal mathematical and logical reasoning benchmarks, including MathVista [36], MathVision [54], MathVerse [68], WeMath [44], DynaMath [70], LogicVista [60], and MMMU [65]. To validate the effectiveness of Athena-ORM and Athena-PRM in the text-only scenario, we conduct experiments on the GSM8K [9] and MATH [18] under the BoN settings. Further details can be found in Appendix A.2.
Settings. We employ Athena-ORM and Athena-PRM as the reward model for BoN evaluation, sampling $N=8$ solutions for every problem by default. For decoding, we set the temperature to 0.8 and use nucleus sampling [20] with top-$p$ set to 0.9, and report the average of five runtimes. For multimodal benchmarks, we choose Qwen2.5-VL-7B/72B [3] and InternVL2.5-8B [8] as policy models. [^4] All models used in this paper are “instruct” version. For simplicity, we ignore “instruct” in the text. For text-only datasets, we choose Qwen2.5-7/72B [61] and Ministral-8B [^5] https://huggingface.co/mistralai/Ministral-8B-Instruct-2410 as policy models. Zero-shot and self-consistency [58] serve as our baselines, alongside a recent public multimodal process reward model, VisualPRM-8B [57].
Training details. To train ORMs and PRMs, we fine-tune Qwen2.5-VL-7B on our dataset as introduced in Sec. 2.3. We use the AdamW optimizer [32] with the weight decay as 0 and set the learning rate as $1e^{-6}$ with cosine decay. We train all models with one epoch and set the global batch size to 64. We use DeepSpeed with zero2 [45] and flash-attention-2 [11] to improve training efficiency. To train ORMs, a new special token <step> is added at the end of solutions for ORMs training. This makes a consistent formulation for ORMs and PRMs. For PRMs, <step> is added to separate all reasoning steps and assign labels to corresponding tokens. We use “$+$” and “--” as labels to denote the correctness of each step. We use 8×8$\times$8 × AMD-MI250 GPUs for training and data synthesis.
Results. Table 1 demonstrates that Athena-ORM and Athena-PRM generally enhance the reasoning performance of MLLMs across policy models of different sizes and benchmarks. Notably, Athena-PRM achieves a +10.2 points improvement on the WeMath dataset with Qwen2.5-VL-7B [3] as the policy model, compared to the zero-shot baseline. ORMs also exhibit general improvement across all benchmarks and policy models, although some gains are limited, such as +0.4 points on MathVerse with Qwen2.5-VL-72B [3]. In contrast, PRMs consistently show substantial performance gains over ORMs, illustrating the benefits of fine-grained rewards in test-time scaling.
Table 2 highlights that Athena-PRM enhances reasoning abilities for text-only inputs in both the Qwen2.5 series and Ministral-8B models. Specifically, Athena-PRM achieves +9.9 points improvement for Ministral-8B on the challenging MATH [18] benchmark, demonstrating its effectiveness in text-only scenarios.
Table 3: Results on VisualProcessBench. We report the F1-score on different dataset and macro F1 in overall. We highlight the best result across all models.
MMMU MathVision MathVerse DynaMath WeMath Overall
Random Guessing 50.0 50.0 50.0 50.0 50.0 50.0
Proprietary Models as Judge
GPT-4o-mini [41] 53.6 58.9 57.1 56.7 58.5 57.9
GPT-4o [41] 56.3 60.2 59.7 59.0 63.3 60.3
Gemini-2.0-Flash [15] 58.5 60.1 62.8 66.7 58.7 62.3
Open-source Models as Judge
MiniCPM-V2.6-8B [62] 44.9 50.9 58.9 46.7 57.4 50.4
LLaVA-OV-7B [26] 45.7 43.0 42.2 44.7 52.5 44.4
LLaVA-OV-72B [26] 46.1 48.4 53.0 57.0 57.3 52.3
Qwen2.5-VL-7B [3] 53.1 51.8 47.8 51.3 54.2 51.0
Qwen2.5-VL-72B [3] 59.2 59.0 59.7 62.9 62.3 60.5
InternVL2.5-8B [8] 47.1 45.5 47.8 50.3 50.8 48.0
InternVL2.5-26B [8] 48.8 47.4 49.2 50.4 51.4 49.2
InternVL2.5-38B [8] 51.5 48.4 50.9 51.8 52.5 50.8
InternVL2.5-78B [8] 52.0 51.7 53.7 50.8 52.5 52.6
Multimodal PRMs
VisualPRM-8B [57] 58.5 62.1 61.0 62.7 61.8 62.0
Athena-PRM 74.1 61.2 60.7 72.7 73.8 65.9
3.2 Evaluation on Direct Judgment of Steps
Setup. In addition to Best-of-N evaluation, we assess Athena-PRM on VisualProcessBench [57], comparing it with general MLLMs and VisualPRM-8B [57]. VisualProcessBench [57] aims to evaluate the ability to directly judge the correctness of each step. For MLLMs, we prompt the model to analyze and judge the correctness of each step. We report the F1-score for every subset and the micro-F1 score overall.
[…]
3.3 Evaluation on the Fine-tuned Model
Table 4: Comparison of multimodal reasoning and mathematical performance across different models. $\Delta$ denotes improvements compared to Qwen2.5-VL-7B [61]. We highlight improvements at least +0.5 points with green.
MMMU WeMath MathVista MathVision MathVerse DynaMath LogicVista
Proprietary Models
GPT-4o [41] 72.9 50.6 71.6 43.8 49.9 48.5 64.4
Claude-3.7-Sonnet [1] 71.0 49.3 66.8 41.9 46.7 39.7 58.2
Gemini-2.0-Pro [15] 72.6 56.5 71.3 48.1 67.3 43.3 53.2
Open-source Models (>70B)
InternVL2.5-78B-MPO [56] 68.2 37.6 76.6 36.2 43.7 21.2 50.8
InternVL2.5-78B [8] 70.1 39.8 70.6 32.2 39.2 19.2 49.0
Qwen2.5-VL-72B [3] 68.2 49.1 74.2 39.3 47.3 35.9 55.7
Open-source Models (~7B)
InternVL2.5-8B [8] 56.2 20.2 58.3 20.0 20.4 9.2 33.6
MiniCPM-o-2.6-8B [62] 50.9 25.2 73.3 21.7 35.0 10.4 36.0
Qwen2.5-VL-7B [3] 58.0 36.2 68.1 25.4 41.1 21.8 47.9
Athena-7B 61.1 43.0 71.4 25.7 45.7 21.9 51.3
$\Delta$ +3.1 +6.8 +3.3 +0.3 +4.6 +0.1 +3.4
Setup. We select Qwen2.5-VL-7B [3] as the initial policy model, obtaining training data as outlined in Sec. 2.4. We use AdamW [32] optimizer and set the learning rate as $1e^{-6}$ with cosine decay. The batch size is set to 128, and the number of epochs is set to one. We denote our fine-tuned model as Athena-7B and evaluate it on the same benchmarks as Sec. 3.1 in the zero-shot setting. For evaluation, we use VLMEvalKit [14] to evaluate Athena-7B and report the results of all baselines from the open leaderboard. [^6] https://huggingface.co/spaces/opencompass/Open_LMM_Reasoning_Leaderboard
Results.   As shown in Table 4, Athena-7B consistently improves upon Qwen2.5-VL-7B [3] across all benchmarks, with significant enhancements of at least $+$0.5 points on five benchmarks. Athena-7B outperforms large open-source models (e.g. InternVL2.5-78B [8]) on five benchmarks. On the MathVista  [36], Athena-7B even achieves the comparable performance with proprietary models such as Gemini-2.0-Pro [15], Claude-3.7-Sonnet [1], and GPT-4o [41].
3.4 Analysis
Test time scaling. To evaluate the scalability of Athena-ORM and Athena-PRM, we compare their performance using different numbers of samples ranging from 4 to 64 on the MathVista [36]. As shown in Figure 2, reward ranking for solutions at test time improves reasoning performance across different policy models and number of samples. Specifically, ORMs and PRMs improve the self-consistency [58] baseline by 0.8 and 2.1 points with Qwen2.5-VL-72B [3] using eight samples, respectively. As the sample size increases, the performance improves steadily. However, the performance gain of ORMs is markedly lower than that of PRMs. For instance, Athena-PRM achieves 82.4% with Qwen2.5-VL-72B [3] using 64 samples, while Athena-ORM achieves similar performance with self-consistency [58] baseline. These findings underscore the advancements and scalability of PRMs in test-time scaling scenarios.
(a) Qwen2.5-VL-7B
(b) InternVL2.5-8B
(c) Qwen2.5-VL-72B
Figure 2: Best-of-N results on the MathVista [36] across different policies. The number of solutions we sample is from 4 to 64.
Table 5: Ablation study of our design choices on the MathVista [54] and WeMath [44] with Qwen2.5-VL-7B [3] as the policy model.
Method Data ORM init Up-sample MathVista WeMath
Zero-shot - - - 68.1 36.2
Self-consistency - - - 71.6 44.7
Athena-ORM Athena-600K - - 72.8 45.1
Athena-PRM Random 5K 73.1 45.2
Athena-5K 74.1 45.8
Random 5K 73.6 45.4
Athena-5K 74.8 46.2
Athena-5K 75.2 46.4
MC-300K 75.5 46.4
The effectiveness of our design choices.  Table 5 presents results under various design choices. Our findings indicate that PRMs trained with selectively chosen 5K samples significantly outperform those trained with randomly selected samples, achieving scores 74.1 vs.73.1 and 74.8 vs.73.6 when initialized from ORMs. Moreover, a large-scale dataset of 300K samples labeled using vanilla MC-based methods yields similar performance to our selected 5K samples. With these selected samples, ORM initialization enhances performance by 0.7 points (74.8 vs.74.1), and up-sampling negative labels further boosts performance by 0.4 points on the MathVista [36].
Computation cost analysis.  Vanilla MC-based estimation [55] incurs substantial computational demands to estimate the correctness of intermediate steps. Our prediction consistency filter strategy significantly reduces this burden, requiring only 5K samples for training, thereby lowering the computational cost by a factor of 60. For data synthesis, we employ vLLM primarily [25] to accelerate MLLM inference. Compared to the MC-300K dataset, our Athena-5K dataset reduces GPU hours by approximately 45-fold while achieving similar performance.
4 Related Work
Test time scaling.  While scaling data and model parameters during training has been extensively studied [23, 19], scaling computation at test time has garnered interest more recently [50, 46]. TTS enables language models to leverage additional computational resources when faced with challenging questions. Our study primarily focuses on parallel scaling [67], wherein language models generate multiple outputs simultaneously and aggregate them into a final answer. This aggregation can involve selecting the most common outcome [58] or using reward models [50] to rank solutions, selecting the one with the highest reward, as demonstrated in our Best-of-N settings.
Process reward models.  Reward models are crucial in reinforcement learning [17, 47, 55, 42, 4, 13] and test time scaling [46, 28, 55, 9, 50]. Outcome reward models [9] assign a scalar value to question-response pairs, whereas process reward models evaluate each step, typically achieving superior performance and generalization [28, 37, 46, 27]. While better performance and generalization are achieved, collecting data to train PRMs is challenging. PRM800K [28] is the first open-source process supervision dataset annotated by humans, but scaling it is challenging due to the time and skill required for annotating reasoning steps. Math-Shepherd [55] proposes an automated method for estimating intermediate step correctness using Monte Carlo estimation, though it generates some incorrect labels and demands extensive computational resources. Our work explores strategies to reduce computational costs while improving process label accuracy.
5 Conclusion
In this paper, we introduce Athena-PRM, trained on 5K high-quality data with process labels. To automatically obtain accurate process labels, we employ a data filtering method based on consistency between weak and strong completers. This strategy significantly reduces computational costs for training and data synthesis while ensuring precise process labels for training PRMs. Additionally, we present two effective strategies for training PRMs: ORM initialization and negative data up-sampling. Athena-PRM demonstrates substantial performance improvements in Best-of-N evaluations, achieving state-of-the-art results on VisualProcessBench. Leveraging Athena-PRM, we train Athena-7B, fine-tuned from Qwen2.5-VL-7B using a reward ranked fine-tuning approach, markedly enhancing the reasoning capabilities of Qwen2.5-VL-7B. Extensive experiments confirm the effectiveness of our proposed methods.
Appendix A More Experiments and Details
A.1 Choices of Completers
Here we verify the effect of combining different completers on the accuracy of estimated labels. Results are listed in Table 6. From Table 6, it is noted that the consistency filter for two completers usually improves the accuracy of estimated labels. For example, combining two weak or strong completers improves accuracy by almost two points in general. Furthermore, the accuracy of estimated labels is improved by combining one weak and one strong completer, achieving the most competitive accuracy, e.g., 95.2% with weak completer $\phi_{w}^{2}$ and strong completer $\phi_{s}^{1}$. This shows that the proposed consistency filter strategy is effective across all combinations of different completers and achieves the best performance when we combine a weak completer and a strong completer.
Table 6: The accuracy of estimated process labels with Different completers combination. In Two Completers, we use consistency filter introduced in Sec. 2.2. $\phi_{w}^{1}$, $\phi_{w}^{2}$, $\phi_{s}^{1}$, and $\phi_{w}^{2}$ denote Mistral-7B-Instruct, Qwen2-7B-Instruct, Qwen2.5-72B-Instruct and Llama-3-70B-Instruct, respectively.
Weak Completer Strong Completer Accuracy(%)
$\phi_{w}^{1}$ $\phi_{w}^{2}$ $\phi_{s}^{1}$ $\phi_{s}^{2}$
Single Completer 78.2
80.1
83.4
84.7
Two Completers 82.3
85.6
94.1
93.8
95.2
94.7
A.2 Benchmarks
We provide more details about used benchmarks in Table 7.
Table 7: Details of all benchmarks used and corresponding metrics.
Split # Samples Metric
MathVista [36] Testmini 1000 Accuracy
MathVision [54] Full 3040 Accuracy
MathVerse [68] Vision-Only 788 Accuracy
DynaMath [70] Full 5050 Accuracy
WeMath [44] Testmini 1740 Score (Strict)
LogicVista [60] Full 448 Accuracy
MMMU [65] Validation 900 Accuracy
GSM8K [9] Test 1319 Accuracy
MATH [18] Test 5000 Accuracy
VisualProcessBench [57] - 2866 F1-score
A.3 Reward Choice
In this paper, we consider three reward choices of Athena-PRM for aggregating step scores into a final score [69], including PRM-minium, PRM-last, and PRM-product. For all reasoning steps $\{a_{i}\}_{i=1}^{K}$, its corresponding reward for each step is $\{r_{i}\}_{i=1}^{K}$, the reward $r$ for each solution with different scoring methods are calculated as follows: 1) PRM-minium uses the minium reward across all steps, i.e., $r=\min\{r_{i}\}_{i=1}^{K}$; 2) PRM-last uses the reward of last step, i.e., $r=r_{K}$; 3) PRM-product uses the reward product of all steps, i.e. $r=\prod_{i=1}^{K}r_{i}$.
We list results in Table 8. It is shown that different choices of selecting a reward for PRMs could produce different performance. In all other experiments, we use the minimum reward across all steps as the reward for the solution unless otherwise specified.
Table 8: Best-of-8 performance comparison on the MathVista [36] with Qwen2.5-VL-7B [3] with different reward choices: last step reward, minium reward across all steps and product of all rewards. Default settings are marked in gray.
Method Accuracy(%)
Zero-shot 68.1
Self-consistency 71.6
ORM 72.8
PRM-product 74.5
PRM-last 75.1
PRM-minium 75.2
A.4 Label Distribution
We present the distribution of process labels in Table 9. It is shown that label imbalance exists in PRM800K and Athena-300K. We try different up-sample rates (e.g. ×$\times 2$) for samples with error steps and find that slightly up-sampling negative data is beneficial for training PRMs, as shown in Table 10. We find that increasing the up-sample rate will not improve performance. We set the up-sample rate to 2 for other experiments.
Table 9: Distribution of labels across different datasets. Good, neutral, and bad denote different process labels.
Dataset Good Neutral Bad
PRM800K [28] 76% 12% 12%
Athena-300K 81% - 19%
Athena-5K 79% - 21%
Table 10: Results on MathVista under Best-of-8 setting with different up-sampling rate. We use Athena-5K as training dataset and Qwen-2.5-VL-7B as the policy model. Default settings are marked in gray.
Method Up-sample rate Accuracy(%)
Zero-shot - 68.1
Self-consistency - 71.6
Athena-ORM - 72.8
Athena-PRM 74.8
Athena-PRM ×$\times 2$ 75.2
Athena-PRM ×$\times 5$ 75.2
Athena-PRM ×$\times 10$ 74.9
Appendix B Limitations and Social Impacts
Limitations.  Athena-PRM achieves impressive performance across various scenarios and benchmarks. However, Athena-PRM is specifically designed to enhance the reasoning capabilities of LLMs and MLLMs. Its application may be limited to solving problems with clear answers, such as multi-step reasoning and math problems. Effectively assigning rewards for intermediate steps in open-ended text generation tasks remains an open problem, and we leave this for future research.
[…]
Social Impacts.   The paper introduces an effective method for generating high-quality, process-labeled data for reasoning problems. Our approach enhances the efficiency and effectiveness of training PRMs. We hope that our research will have a broader impact on improving the reasoning abilities of foundation models. Since our method focuses on general improvements in the reasoning capabilities of foundation models, it does not directly produce negative societal impacts.
Appendix C Case Study
We provide some examples from VisualProcessBench and MathVista to help readers understand how PRMs work, see Figure 3456.
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: When the ant <image1> walks from home <image2> along the arrows 3,3,3,1$\rightarrow$ 3$\uparrow$ 3$\rightarrow$ 3$\uparrow$ 1→ 3 , ↑ 3 , → 3 , ↑ 1, he gets to the ladybird <image3>. Which animal does the ant <image1> get to when he walks from home <image2> along the following arrows: 2,2,3,3,2,2$\rightarrow$ 2$\downarrow$ 2$\rightarrow$ 3$\uparrow$ 3$\rightarrow$ 2$\uparrow$ 2→ 2 , ↓ 2 , → 3 , ↑ 3 , → 2 , ↑ 2? <image6> <image7> Choices:(A) A (B) B (C) C (D) D (E) E (Answer: A)
Step Reward Label
To determine which animal the ant reaches, we will follow the given movements step-by-step on the grid: 0.95
1. Start at the home position (‘<image2>‘), which is at the bottom-left corner of the grid. 0.99
2. Follow the arrows: - $\rightarrow 2$: Move 2 squares to the right. - $\downarrow 2$: Move 2 squares down. - $\rightarrow 3$: Move 3 squares to the right. - $\uparrow 3$: Move 3 squares up. - $\rightarrow 2$: Move 2 squares to the right. - $\uparrow 2$: Move 2 squares up. 0.99
3. Trace the movement: - Start at the bottom-left corner. - Move 2 squares right: You are now in the third column of the first row. - Move 2 squares down: You are now in the third column of the third row. - Move 3 squares right: You are now in the sixth column of the third row. - Move 3 squares up: You are now in the sixth column of the top row. - Move 2 squares right: You are now in the eighth column of the top row. - Move 2 squares up: You are now in the eighth column of the top row (still at the butterfly position). 0.12
The ant reaches the butterfly, which corresponds to option A. 0.87
### Final Answer: A 0.75
Figure 3: A case study from VisualProcessBench [57].
According to the question shown in the image, please directly answer the question and provide the correct option letter, e.g., A, B, C, D. (Answer: A)
Step Reward Label
To find the degree of angle A, we can use the property that the angle subtended by an arc at the center of the circle is twice the angle subtended at any point on the circle. 0.82
1. Identify the key points:
- The arcs are AC and BC.
- Angle A is subtended by these arcs at point O (the center of the circle).
0.85
2. Understand the given information:
- CD is a tangent at point C.
- The angle between the tangent and the chord is equal to the angle subtended by the chord at the circumference. Therefore, angle CDB = 50°.
0.85
3. Angle properties:
- The angle A is subtended by arc BC at the center.
- Since CD is a tangent, angle A = 50° (as the inscribed angle is half the central angle).
0.23
Thus, the degree of angle A is 20°. 0.67
Final answer: A 0.96
Figure 4: A case study from VisualProcessBench [57].
Hint: Please answer the question requiring an integer answer and provide the final value, e.g., 1, 2, 3, at the end. Question: what is the highest number shown on the black outer part of the watch? (Answer: 55)
Step Reward Label
To determine the highest number shown on the black outer part of the watch, let’s analyze the image step by step 0.90
1. The black outer part of the watch appears to be a bezel with numbers printed on it. 0.87
2. The numbers on the bezel are arranged in a circular pattern. 0.96
3. Observing the numbers, they range from 0 to 60, increasing in increments of 5. 0.83
4. The highest number visible on the bezel is 60. 0.34
Therefore, the highest number shown on the black outer part of the watch is 60. 0.15
[…]
Figure 5: A case study from MathVista [36] with Qwen2.5-VL-72B [3].
Hint: Please answer the question requiring an integer answer and provide the final value, e.g., 1, 2, 3, at the end. Question: Subtract all large green matte cubes. Subtract all big green blocks. How many objects are left? (Answer: 6)
Step Reward Label
To solve this problem, let’s identify and count the objects in the image: 0.95
1. There is one large green matte cube. 0.99
2. There are no other big green blocks since the only green block is already counted. 0.90
Now, we subtract these objects as per the instructions 0.92
- Subtract all large green matte cubes: This removes the large green matte cube from the count. 0.99
- Subtract all big green blocks: Since there are no other big green blocks, nothing additional is removed. 0.99
After removing the large green matte cube, the remaining objects are: - A small gray metallic cylinder. - A large gold metallic sphere. - A small gray matte sphere. - A large cyan matte cylinder. - A small cyan matte sphere. - A large gold metallic cube. 0.84
Counting the remaining objects, we have 6 objects left. 0.92
[…]
Figure 6: A case study from MathVista [36] with Qwen2.5-VL-72B [3].